Thanks for the link. It seems the LPF selection works fine, as voltages are like in the document also for the bands without any significant output power.
However it seems that even though the clock2 signal from IC204 looks fine:
But after IC501 and IC502 it does not look fine anymore
and after IC503 it looks even worse (at BS170 gates):
I suppose IC501 and IC502 ought to be replaced first?
(O-scope probe in 10x mode in all screenshots above, QMX+ in 40m CW mode)
Regards, Jonas - SM4VEY
"
PS. The list of broken and replaced components after the calibration "event" so far extends?to Q108, Q112, Q106, D104, D108, L101, Q503, Q504, Q505 and Q506. Soon to be adding IC501 and IC502 as well. I suppose it would have been better?to trash this one and just move the Toroids over to another board... oh well, hindsight is 20/20 :-)? DS.
Den m?n 19 maj 2025 kl 22:09 skrev DH8WN via <DH8WN=[email protected]>: